OpenStudy (anonymous):

Simplify 125 to the power of negative 2 over 3

jimthompson5910 (jim_thompson5910):

\[\Large 125^{-\frac{2}{3}}\] \[\Large \frac{1}{125^{\frac{2}{3}}}\] \[\Large \frac{1}{\left(\sqrt[3]{125}\right)^{2}}\] \[\Large \frac{1}{\left(\sqrt[3]{5^3}\right)^{2}}\] \[\Large \frac{1}{\left(5\right)^{2}}\] \[\Large \frac{1}{25}\] So \[\Large 125^{-\frac{2}{3}}=\frac{1}{25}\]